### v2

@@ B.9 Proof of Proposition 5
$\leq\sum_{i\neq 1}\Pr\left(|D^{i}_{k}|\geq\frac{\varepsilon}{\sigma_{k}}\right)\xrightarrow[\rho\to 1]{}0.$ (118)
@@ B.11 Random Favorite (RF) Mechanism
Accordingly, given $(a,b)$, the demand for object $1$ is $a^{2}/(a+b)$ and the demand for object $2$ is $b^{2}/(a+b)$. The market clearing condition requires that $a^{2}/(a+b)=m_{1}$ and $b^{2}/(a+b)=m_{2}$. Solving these equations, we obtain $a=m_{1}+\sqrt{m_{1}m_{2}}$ and $b=m_{2}+\sqrt{m_{1}m_{2}}$ as a unique $(a,b)$ satisfying the agent’s optimality condition and the market clearing condition.
The residual surplus achieved by RF is
@@
Since we assume $m_{1}\geq m_{2}$, $RS(\mathcal{M}_{RF})\geq RS(\mathcal{M}_{SD})$ always holds, and the equality holds if and only if $m_{1}=m_{2}$. In Figure A.2, we plot

### v3

@@ B.9 Proof of Proposition 5
$=\sum_{i\neq 1}\Pr\left(|D^{i}_{k}|\geq\frac{\varepsilon}{\sigma_{k}}\right)\xrightarrow[\rho\to 1]{}0.$ (124)
@@ B.11 Random Favorite (RF) Mechanism
Accordingly, given $(a,b)$, the demand for object $1$ is $a^{2}/(a+b)$ and the demand for object $2$ is $b^{2}/(a+b)$. The market clearing condition requires that $a^{2}/(a+b)=m_{1}$ and $b^{2}/(a+b)=m_{2}$. Solving these equations, we obtain
$a=m_{1}+\sqrt{m_{1}m_{2}},\quad b=m_{2}+\sqrt{m_{1}m_{2}},$ (136)
which constitute the unique solution in $(0,1]^{2}$ satisfying the agent’s optimality condition and the market-clearing condition whenever $m_{1}+\sqrt{m_{1}m_{2}}\leq 1$. We focus on the case in which such an interior solution exists.
In this interior region, the residual surplus achieved by RF is given by
@@
Since we assume $m_{1}\geq m_{2}$, within the interior region, $RS(\mathcal{M}_{RF})\geq RS(\mathcal{M}_{SD})$ always holds, and the equality holds if and only if $m_{1}=m_{2}$. In Figure A.2, we plot